OpenStudy (anonymous):

Ram's present age is three times his son's present age and twofifth of his father's present age. The average of the present ages of all of them is 46 years . Wha t i s the difference between the R a m ' s son' s present age and Ram's father's present age ?

OpenStudy (anonymous):

too much

OpenStudy (anonymous):

Three equations and three variables. solve for them.

OpenStudy (anonymous):

r=3s

OpenStudy (anonymous):

thank u very much

OpenStudy (anonymous):

\[ r = 3s \\ r = 2f/5\\ (s+r+f)/3 = 46 \]

OpenStudy (anonymous):

s=r/3,f=5r/2 so 23r/6=46. r=12 so12-36=24

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

the answer is 39